# references.bib

### v3

@@ 4.1 Zero-Shot Inference Is A Strong FSOD Baseline
We compare state-of-the-art FSOD methods with zero-shot inference from GroundingDINO [liu2023grounding] on COCO in Table 1. Surprisingly, GroundingDINO outperforms DiGeo [ma2023digeo] by 16.2% AP averaged across both base and novel categories despite never being trained on COCO images. GroundingDINO’s impressive performance is due to its large-scale multi-modal pre-training on Objects365 [shao2019objects365], GoldG [goldg] and Cap4M [li2021grounded]. It is worth noting that GroundingDINO achieves higher AP on novel classes than base, suggesting that novel classes in existing benchmarks (e.g. car and person) are actually not rare in the real world. Therefore, FSOD benchmarks should be re-framed to reflect real-world applications, motivating our setup.
@@ 4.2 Foundational FSOD with nuImages
In the context of foundational models, we argue that partitioning datasets into base and novel classes no longer makes sense. Instead, FSOD methods should only train on $K$-shot annotations for $C$ target classes, and also evaluate performance on these $C$ classes. We pre-train TFA [wang2020frustratingly] on diverse datasets and fine-tune on $K$ examples per class and highlight model performance in Table 2. We train two variants of TFA trained on COCO-base and LVIS-base and fine-tune both models on $K$ examples of the nuImages classes. Surprisingly, both variants of TFA achieve less than $3$ AP (cf. Table 2). We posit that this is largely due to poor classification performance. Since both LVIS and COCO classes do not significantly overlap with nuImages classes, learning a classifier from few examples is extremely difficult. However, we find that simply re-training TFA with a frozen CLIP-based classifier (similar to Detic) dramatically increases performance, reiterating the utility of language and web-scale pre-training in data-constrained settings.
Table 3: Empirical Analysis of Baselines (10-Shots) on our Benchmark. We evaluate popular VLMs on the nuImages FSOD Benchmark and find that MQ-GLIP performs the best among all baseline models. Notably, it achieves 17.0 mAP zero-shot language-only performance, and achieves 21.4 mAP via zero-shot multi-modal prompting averaged over all classes. Remarkably, our 2024 competition winners further improved performance to 45.4 mAP, beating our best baseline by 24.0%.
@@ 4.3 Empirical Analysis of Results
We evaluate several popular VLMs on the nuImages Foundational FSOD (10-shots) benchmark and present salient insights from Table 3 below.
Zero-Shot Detection. Somewhat unsurprisingly, we find that (i) greater pre-training data scale and diversity, along with (ii) larger backbones result in better zero-shot performance. Notably, GLIP achieves 17.01% zero-shot performance, surpassing all other methods trained with less data and smaller backbones.
Prompt Engineering. We attempt to improve zero-shot performance using synonyms for class names derived from the annotator textual instructions. We see minor improvements (e.g. Detic improves from $14.40$ mAP $\rightarrow 14.92$ mAP), indicating that leveraging rich textual descriptions beyond class names can improve concept alignment.
@@
Multi-Modal Prompting. We observe that Multi-Modal Prompting (MQ-GLIP) achieves the best performance ($21.42$ mAP) out of all open-source methods in Table 3. We attribute this to its large pre-trained dataset, bigger backbone (SWIN-L) and multi-modal prompts used during inference. Notably, the benefit of multi-modal prompts can be seen by comparing MQ-GLIP ($21.42$ mAP) against MQ-GLIP-Image ($14.07$ mAP), which uses visual prompting and MQ-GLIP-Text ($17.01$ mAP), which uses language prompting. Interestingly, MQ-GLIP does not require gradient-based fine-tuning, which differs from all existing conventional few-shot methods. Therefore, we posit that future few-shot methods should further explore in-context learning. Just as multi-modal annotator instructions aid human annotator alignment, we find that multi-modal prompting significantly improves VLM concept alignment.
Multi-Modal Chat Agents. Given the strong performance of GPT-4o for general visual understanding, we repurpose it for our task by prompting the model to re-classify image patches from Detic’s RPN. Specifically, we ask GPT-4o to predict a class and confidence for each image crop. Surprisingly, we observe reasonable performance (9.95 mAP) despite GPT-4o not being trained as an object detector, emphasizing the importance of the scale of pre-training data. We explore iterative prediction refinement in the supplement.
CVPR 2024 Challenge. Finally, we highlight our top three submissions (out of six participants) from the inaugural Foundational FSOD challenge. Notably all top performers beat our baselines, with the winning team achieving $45.35$ AP! We discuss more details in the supplement.
@@
Table 4: Random Split vs “Best” Split. We construct the “best" split by selecting per-class few-shot examples that lead to the highest performance on a held-out set. Unsurprisingly, the best split performs better than any random split, especially for very limited data settings (e.g. $5$-shot detection). This evaluation setting closely mimics how human annotators are “aligned” to target concepts, since annotator guides are constructed using hand-picked iconic visual examples.
@@ 4.5 Limitations and Future Work
Despite using VLMs pre-trained on large-scale datasets, we find that performance for rare categories (defined by the cardinality of each class in the original dataset) is considerably lower than for common classes. We posit that VLMs are pre-trained with imbalanced data which includes many examples of common categories like truck but few examples of rare categories like stroller. Our work does not explicitly improve detection performance on rare classes. Interestingly, since VLMs like Detic [zhou2022detecting], GLIP [li2021grounded], and GroundingDINO [liu2023grounding] are trained with different data sources, each model has dramatically different zero-shot performance on novel categories like stroller. Ensembling predictions from different VLMs may yield better detection accuracy for rare categories. In addition, although our work motivates the use of rich textual descriptions found in instructions for multi-modal alignment, our current results use only nouns (class names and synonyms) as text prompts.
@@
Out-of-Domain Benchmarks. Another way to address benchmarking is to collect test scenarios that are designed to be dissimilar from internet images. For example, out-of-domain images may include medical data (though foundational performance is still surprisingly effective [Wang2022MedCLIPCL]). We admittedly did not take this route, since urban imagery is similar to images found online and arguably many applications of interest fall under this category. Moreover, there exist ample opportunity for technical innovation in this setting (as suggested by our CVPR 2024 challenge results!). Alternatively, one can manually collect and sequester images that will never be released on the internet. Since ensuring privacy may itself be challenging, yet another approach is to leverage the continual learning paradigm, where new test sets are continually constructed over time.
Comparing Models. Fairly comparing foundation models requires careful consideration. Although accuracy is a valuable metric, it is intrinsically tied to the scale of pre-training data and model architecture. Notably, the LLM community already ranks models via a Pareto frontier of accuracy vs. parameter count. We advocate for a similar approach for Foundational FSOD that considers backbone architecture (e.g. ResNet-50 vs. Swin-B) and pre-training datasets (e.g. CC4M, GoldG, LVIS).
@@ Appendix A Baseline Implementation Details
Appendix B CVPR 2024 Competition Details.
Table 6: CVPR 2024 Foundational FSOD Competition Results.
Team Name Average Precision (AP)
All Many Medium Few
PHP_hhh 45.35 64.25 53.43 20.19
NJUST KMG 32.56 50.21 34.87 15.16
zjyd_cxy_vision 31.57 46.59 33.32 17.03
Baseline (MQ-GLIP) 21.51 32.25 23.35 10.41
team_anon 17.36 25.29 21.93 5.42
youyouqiu 13.16 11.29 19.20 7.68
zhao 11.38 11.16 16.76 5.30
zjdcxy 7.80 5.44 13.43 3.20
Our inaugural Foundational FSOD competition (hosted on Eval AI) received submissions from eight teams (some submissions are private) around the world. We present a ranked list of participants at the close of our competition on June 7th AOE in Table 6. Notably, three teams were able to beat our MQ-GLIP baseline. Unfortunately, the top performing team wasn’t willing to publicly share details about their method. We summarize contributions from the other two top teams below.
NJUST KMG presents a method leveraging a vision-language model (VLM) enhanced with a multimodal large language model (MM-LLM) to improve Few-Shot Object Detection (FSOD). To address the challenge of misalignment between VLMs and target concepts, authors propose generating descriptive referential expressions for each category using MM-LLM. This involves annotating images with bounding boxes, prompting ChatGPT to provide descriptive terms for each object, and then creating multiple referential expressions by randomly combining these terms. The VLMs then select the best referential expression for each category by matching the maximum Intersection over Union (IoU) in the training set, and these expressions are used to generate pseudo-labels for all training images, which are combined with original labeled data to fine-tune the VLM. This iterative process of pseudo-label generation and optimization significantly enhances the VLM’s performance.
ZJYD CXY Vision proposes Instruction DINO (ISD), a DETR-based detector architecture and incorporates early fusion of image and text information, using a Swin-L visual backbone and EVA02-CLIP-L text encoder. Pre-training involves two stages using various datasets, transforming grounding data into single-object descriptions with QWen Max. For few-shot fine-tuning, the model adopts a flexible training format and uses VLMs like CLIP, TAP, and LLava for negative sample generation, finding that prompt tuning and text encoder fine-tuning generalize better than visual encoder fine-tuning. The final fine-tuning method combines prompt tuning and negative sampling, significantly improving mAP. To address sparse annotations, the visual encoder is initially fine-tuned to generate pseudo-label annotations, which are then used to complete training with prompt tuning.
Figure 5: Iteratively Prompting ChatGPT. Despite its large-scale pre-training, multi-modal models like ChatGPT-4o also suffers from concept alignment. Specifically, GPT-4o makes highly confident but incorrect predictions for debris. We propose an iterative prompting strategy to better align the model to a target concept. Given a few visual examples per-class from the training-set, we query ChatGPT to use its “web-scale knowledge” to generate text descriptions. We then augment the input to MQDet to incorporate this additional context for zero-shot evaluation.
Appendix C Iterative Prompting with Multi-Modal Chat Assistants
Typically, clients provide human annotators with a set of multi-modal instructions and a corpus of unlabeled data for annotation. Annotators start by first labeling a small subset of the data for review by the client, who acts as a domain expert and provides feedback on erroneous annotations (highlighting concept misalignment). Annotators use this feedback to annotate another subset of the data. This iterative process continues until the client is satisfied with the annotators’ ability to accurately label the entire dataset.
As shown in Figure 5, we explore the idea of iteratively prompting multi-modal chat assistants like ChatGPT to mimic the real-world workflow of human annotators. We start by asking GPT-4o to classify image crops of debris (derived from the few-shot training split). Notably, GPT-4o incorrectly classifies these training examples with high confidence. Therefore, we prompt GPT-4o to generate its own text descriptions of the few-shot examples according to its “web-scale knowledge”. Finally, we use the class names, generated text descriptions for debris, and few-shot visual examples to MQDet to predict instances of debris in the test-set.
We find that prompting MQDet with class names, ChatGPT generated text descriptions, and few-shot visual examples improves performance by 0.67% ($21.42$ mAP $\rightarrow$ $22.09$ mAP) over the baseline. Interestingly, although debris does not change when prompted with generated text descriptions, pushable pullable ($3.6$ AP $\rightarrow$ $15.29$ AP) and barrier ($11.6$ AP $\rightarrow$ $15.31$ AP) accuracy improve significantly. We posit that this improvement is due to the reduction in confusion (or the over-confident incorrect predictions) between debris and pushable-pullable (and barrier). Surprisingly, one of the top submissions to our CVPR challenge also use ChatGPT to generate meaningful text descriptions to improve detection concept alignment.
Appendix D Analysis of Federated Fine-Tuning
@@
Appendix E Impact of Box-Level Supervision for Foundational FSOD
We evaluate the importance of using bounding-box supervised data in pre-training. Unlike Detic, which trains on box-supervised data from LVIS, COCO and image-text data from ImageNet21k, RegionCLIP[zhong2022regionclip] only pre-trains on image-text pairs from the Conceptual Captions (CC3M) dataset [cc4m].
We report RegionCLIP’s zero-shot and fine-tuning performance on nuImages averaged over $3$ random splits in Table 8. Detic zero-shot outperforms RegionCLIP zero-shot by $\sim 12$ AP ($14.26$ vs $2.34$). While fine-tuning RegionCLIP improves overall performance, Detic achieves higher accuracy for $K=\{5,10,30\}$ shots. This highlights the importance of supervision type (e.g. box-supervised data) and data scale used for pre-training.
@@ Appendix E Impact of Box-Level Supervision for Foundational FSOD
Appendix F NuImages Annotator Instructions
@@
Table 10: Empirical Analysis of Baselines (5-Shots) on nuImages.
@@
Table 11: Empirical Analysis of Baselines (30-Shots) on nuImages.
@@
Appendix G Empirical Analysis of Baselines (5-Shots and 30-Shots)
We evaluate all baselines for the nuImages experiments with 5-shots and 30-shots in Tables  10 and 3 respectively. We find that trends from the main paper hold. Notably, MQ-GLIP with-multi-modal prompting performs the best. However, we find that adding more examples (e.g. MQ-GLIP 5-shot vs. MQ-GLIP 30-shot) doesn’t seem to help in-context learning based methods nearly as much as gradient-based fine-tuning approaches.
Appendix H Foundational FSOD with LVIS
Table 12: LVIS Foundational FSOD Performance. We present fine-tuning results for different variants of Detic on the LVIS 10-shot dataset. We follow the standard FSOD setup and pre-train Detic on LVIS-base for fair comparison with prior work. Detic pre-trained only on LVIS-base outperforms specialized methods like TFA and DiGeo by $\sim$6 AP, without fine-tuning on rare classes. Since we keep the model backbone (ResNet-50) and pre-training data same for all methods, these performance improvements can be attributed to Detic’s CLIP-based classifier. This demonstrates that concept leakage through language significantly improve FSOD, and leveraging language cues should be embraced in data constrained settings. Naively fine-tuning Detic yields a performance drop of $AP_{f}$ and $AP_{c}$ because treating common classes as negatives in rare category federated datasets hurts performance. Instead, we find that embracing the federated nature of FSOD datasets provides consistent improvements in fine-tuning (30.0 vs. 30.8 for ResNet-50). Further, pseudo-labeling negatives in each image provides a modest improvement (30.8 vs. 31.6 for ResNet-50). Similar trends hold for the Swin backbone.

### v4

@@ 4.1 Zero-Shot Inference Is A Strong FSOD Baseline
We compare state-of-the-art FSOD methods with zero-shot inference from GroundingDINO [liu2023grounding] on COCO in Table 1. Surprisingly, GroundingDINO outperforms DiGeo [ma2023digeo] by 16.2% AP averaged across both base and novel categories despite never being trained on COCO images. GroundingDINO’s impressive performance is due to its large-scale multi-modal pre-training on Objects365 [shao2019objects365], GoldG [goldg] and Cap4M [li2021grounded]. It is worth noting that GroundingDINO achieves higher AP on novel classes than base, suggesting that novel classes in existing benchmarks (e.g., car and person) are actually not rare in the real world. Therefore, FSOD benchmarks should be re-framed to reflect real-world applications, motivating our setup.
@@ 4.2 Foundational FSOD with nuImages
In the context of foundational models, we argue that partitioning datasets into base and novel classes no longer makes sense. Instead, FSOD methods should only fine-tune on $K$-shot annotations for $C$ target classes, and also evaluate performance on these $C$ classes. We pre-train TFA [wang2020frustratingly] on diverse datasets and fine-tune on $K$ examples per class and highlight model performance in Table 2. We train two variants of TFA trained on COCO-base and LVIS-base and fine-tune both models on $K$ examples of the nuImages classes. Surprisingly, both variants of TFA achieve less than $3$ AP (cf. Table 2). We posit that this is largely due to poor classification performance. Since both LVIS and COCO classes do not significantly overlap with nuImages classes, learning a classifier from few examples is extremely difficult. However, we find that simply re-training TFA with a frozen CLIP-based classifier (similar to Detic) dramatically increases performance, reiterating the utility of language and web-scale pre-training in data-constrained settings.
Table 3: Empirical Analysis of Baselines (10-Shot) on our Benchmark. We evaluate popular VLMs on the nuImages FSOD Benchmark and find that MQ-GLIP performs the best among all baseline models. Notably, it achieves 17.0 mAP zero-shot language-only performance, and achieves 21.4 mAP via zero-shot multi-modal prompting averaged over all classes. We can iteratively prompt GPT-4o for synonyms to describe each of the few-shot examples to expand MQ GLIP’s text prompts, further improving performance by 0.6%. Remarkably, our 2024 competition winners further improved performance to 45.4 mAP, beating our best baseline by 23.3%.
@@ 4.3 Empirical Analysis of Results
We evaluate several popular VLMs on the nuImages Foundational FSOD (10-shot) benchmark and present salient insights from Table 3 below.
Zero-Shot Detection. Somewhat unsurprisingly, we find that (1) greater pre-training data scale and diversity, along with (2) larger backbones result in better zero-shot performance. Notably, GLIP achieves 17.01% zero-shot performance, surpassing all other methods trained with less data and smaller backbones.
Prompt Engineering. We attempt to improve zero-shot performance using synonyms for class names derived from the annotator textual instructions. We see minor improvements (e.g., Detic improves from $14.40$ mAP $\rightarrow 14.92$ mAP), indicating that leveraging rich textual descriptions beyond class names can improve concept alignment.
@@
Multi-Modal Prompting. We observe that Multi-Modal Prompting (MQ-GLIP) achieves the best performance ($21.42$ mAP) out of all open-source methods in Table 3. We attribute this to its large pre-trained dataset, bigger backbone (SWIN-L) and multi-modal prompts used during inference. Notably, the benefit of multi-modal prompts can be seen by comparing MQ-GLIP ($21.42$ mAP) against MQ-GLIP-Image ($14.07$ mAP), which uses visual prompting and MQ-GLIP-Text ($17.01$ mAP), which uses language prompting. Interestingly, MQ-GLIP does not require gradient-based fine-tuning, which differs from all existing conventional few-shot methods. Therefore, we posit that future few-shot methods should further explore in-context learning. Just as multi-modal examples aid human annotator alignment, multi-modal prompting significantly improves VLM concept alignment.
Figure 3: Iteratively Prompting ChatGPT. Despite its large-scale pre-training, multi-modal models like ChatGPT-4o also suffers from concept misalignment. Specifically, GPT-4o makes highly confident but incorrect predictions for debris. We propose an iterative prompting strategy to better align the model to a target concept. Given a few visual examples per-class from the training-set, we query ChatGPT to use its “web-scale knowledge” to generate text descriptions. We then augment the input to MQDet to incorporate this additional context for zero-shot evaluation.
Multi-Modal Chat Agents. As shown in Figure 3, we explore the idea of iteratively prompting multi-modal chat assistants like ChatGPT to mimic the real-world workflow of human annotators. Given the strong performance of GPT-4o for general visual understanding, we repurpose it for our task by prompting the model to re-classify image patches from Detic’s RPN. Specifically, we ask GPT-4o to predict a class and confidence for each image crop. Surprisingly, we observe reasonable performance (9.95 mAP). However, we find that GPT-4o often incorrectly classifies many image crops with high confidence. Therefore, we prompt GPT-4o to generate its own text descriptions of the few-shot examples according to its “web-scale knowledge”. Finally, we use the class names, generated text descriptions, and few-shot visual examples to prompt MQDet to predict new instances of target classes in the test-set. We find that expanding MQDet’s in-context prompt with class names, ChatGPT generated text descriptions, and few-shot visual examples improves performance by 0.67% ($21.42$ mAP $\rightarrow$ $22.09$ mAP) over the MQ-GLIP baseline. Interestingly, although debris accuracy does not change when prompted with generated text descriptions, pushable pullable ($3.6$ AP $\rightarrow$ $15.29$ AP) and barrier ($11.6$ AP $\rightarrow$ $15.31$ AP) accuracy improve significantly. We posit that this improvement is due to the reduction in confusion (or over-confident incorrect predictions) between debris and pushable-pullable (and barrier). Surprisingly, a top submission to our CVPR challenge also used ChatGPT to generate meaningful text descriptions to improve concept alignment.
CVPR 2024 Challenge. Our inaugural Foundational FSOD competition (hosted on Eval AI) received submissions from seven teams (some submissions are private). We present a ranked list of participants at the close of our competition on June 7th 2024 AOE in Table 4. Notably, three teams beat our baselines, with the winning team achieving $45.35$ AP! Unfortunately, the top performing team was not willing to publicly share details about their method. We summarize contributions from the other two top teams below.
Table 4: CVPR 2024 Foundational FSOD Competition Results.
Team Name Average Precision (AP)
All Many Medium Few
PHP_hhh 45.35 64.25 53.43 20.19
NJUST KMG 32.56 50.21 34.87 15.16
zjyd_cxy_vision 31.57 46.59 33.32 17.03
Baseline (MQ-GLIP) 21.51 32.25 23.35 10.41
team_anon 17.36 25.29 21.93 5.42
youyouqiu 13.16 11.29 19.20 7.68
zhao 11.38 11.16 16.76 5.30
zjdcxy 7.80 5.44 13.43 3.20
NJUST KMG presents a method that leverages both VLMs and multi-modal chat agents for Foundational FSOD. To address the challenge of misalignment between GroundingDINO and the target concepts of interest, authors generate descriptive referential expressions by prompting ChatGPT to provide descriptive terms for each few-shot instance. The best referential expression for each category is selected by maximizing the Intersection over Union (IoU) between predictions and the ground truth in the few-shot training set. These referential expressions are then used to generate pseudo-labels for all training images. Lastly, GroundingDINO is fine-tuned on a combination of pseudo-labels and ground-truth instances. The full technical report is available here.
ZJYD CXY Vision proposes Instruction DINO (ISD), a DETR-based detector architecture that incorporates early fusion of image and text information using a Swin-L visual backbone and EVA02-CLIP-L text encoder. Authors use VLMs like CLIP, TAP, and Llava for negative sample generation (similar to our Federated Fine-Tuning). Authors find that prompt tuning and text encoder fine-tuning generalize better than visual encoder fine-tuning. Similar to NJUST KMG, authors first generate pseudo-label annotations for unlabeled categories before fine-tuning on a combination of pseudo-labels and ground truth instances. The final method combines prompt tuning and negative sampling, significantly improving mAP. The full technical report is available here.
@@
Table 5: Random Split vs “Best” Split. We construct the “best" split by selecting per-class few-shot examples that lead to the highest performance on a held-out set. Unsurprisingly, the best split performs better than any random split, especially for very limited data settings (e.g., $5$-shot detection). This evaluation setting closely mimics how human annotators are “aligned” to target concepts, since annotator guides are constructed using hand-picked iconic visual examples.
@@ 4.5 Limitations and Future Work
Despite using VLMs pre-trained on large-scale datasets, we find that performance for rare categories (defined by the cardinality of each class in the original dataset) is considerably lower than for common classes. We posit that VLMs are pre-trained with imbalanced data which includes many examples of common categories like truck but few examples of rare categories like stroller [parashar2024neglected]. Our work does not explicitly improve detection performance on rare classes. Interestingly, since VLMs like Detic [zhou2022detecting], GLIP [li2021grounded], and GroundingDINO [liu2023grounding] are trained with different data sources, each model has dramatically different zero-shot performance on novel categories like stroller. Ensembling predictions from different VLMs may yield better detection accuracy for rare categories. In addition, although our work motivates the use of rich textual descriptions found in instructions for multi-modal alignment, our current results use only nouns (class names and synonyms) as text prompts.
@@
Out-of-Domain Benchmarks. Another way to address benchmarking is to collect test scenarios that are designed to be dissimilar from internet images. For example, out-of-domain images may include medical data (though foundational performance is still surprisingly effective [Wang2022MedCLIPCL]). We admittedly did not take this route, since urban imagery is similar to images found online and arguably many applications of interest fall under this category. Moreover, there exist ample opportunity for technical innovation in this setting (as suggested by our CVPR 2024 challenge results!). Alternatively, one can manually collect and sequester images that will never be released on the internet. Since ensuring privacy may itself be challenging, yet another approach is to leverage the continual learning paradigm [lin2022continual], where new test sets are continually constructed over time.
Comparing Models. Fairly comparing foundation models requires careful consideration. Although accuracy is a valuable metric, it is intrinsically tied to the scale of pre-training data and model architecture. Notably, the LLM community already ranks models via a Pareto frontier of accuracy vs. parameter count. We advocate for a similar approach for Foundational FSOD that considers backbone architecture (e.g., ResNet-50 vs. Swin-B) and pre-training datasets (e.g., CC4M, GoldG, LVIS).
@@ 5 Conclusion
6 Acknowledgements
This research was supported by Bosch Center for AI (BCAI). We thank BCAI for their financial support and resources, which made this work possible. This work was also supported in part by the Institute of Collaborative Innotation and the University of Macau (SRG2023-00044-FST) and the NSF GRFP (Grant No. DGE2140739).
@@ Appendix A Baseline Implementation Details
Appendix B Analysis of Federated Fine-Tuning
@@
Appendix C Impact of Box-Level Supervision for Foundational FSOD
We evaluate the importance of using bounding-box supervised data in pre-training. Unlike Detic, which trains on box-supervised data from LVIS, COCO and image-text data from ImageNet21k, RegionCLIP[zhong2022regionclip] only pre-trains on image-text pairs from the Conceptual Captions (CC3M) dataset [cc4m]. We report RegionCLIP’s zero-shot and fine-tuning performance on nuImages averaged over $3$ random splits in Table 8. Detic zero-shot outperforms RegionCLIP zero-shot by $\sim 12$ AP ($14.26$ vs $2.34$). While fine-tuning RegionCLIP improves overall performance, Detic achieves higher accuracy for $K=\{5,10,30\}$ shots. This highlights the importance of supervision type (e.g. box-supervised data) and data scale used for pre-training.
@@ Appendix C Impact of Box-Level Supervision for Foundational FSOD
Appendix D NuImages Annotator Instructions
@@
Table 10: Empirical Analysis of Baselines (5-shot) on nuImages.
@@
Table 11: Empirical Analysis of Baselines (30-shot) on nuImages.
@@
Appendix E Empirical Analysis of Baselines (5-shot and 30-shot)
We evaluate all baselines for the nuImages experiments with 5-shot and 30-shot in Table 10 and 11, respectively. We find that trends from the main paper hold. Notably, MQ-GLIP with-multi-modal prompting performs the best. However, we find that adding more examples (e.g. MQ-GLIP 5-shot vs. MQ-GLIP 30-shot) does not seem to help in-context learning based methods nearly as much as gradient-based fine-tuning approaches.
Appendix F Foundational FSOD with LVIS
Table 12: LVIS Foundational FSOD Performance. We present fine-tuning results for different variants of Detic on the LVIS 10-shot dataset. We follow the standard FSOD setup and pre-train Detic on LVIS-base for fair comparison with prior work. Detic pre-trained only on LVIS-base outperforms specialized methods like TFA and DiGeo by $\sim$6 AP, without fine-tuning on rare classes. Since we keep the model backbone (ResNet-50) and pre-training data same for all methods, these performance improvements can be attributed to Detic’s CLIP-based classifier. This demonstrates that concept leakage through language significantly improve FSOD, and leveraging language cues should be embraced in data constrained settings. Naively fine-tuning Detic yields a performance drop of $AP_{f}$ and $AP_{c}$ because treating common classes as negatives in rare category federated datasets hurts performance. Instead, we find that embracing the federated nature of FSOD datasets provides consistent improvements in fine-tuning (30.0 vs. 30.8 for ResNet-50). Further, pseudo-labeling negatives in each image provides a modest improvement (30.8 vs. 31.6 for ResNet-50). Similar trends hold for the SWIN-B and SWIN-L backbones.